A particle is dropped vertically from rest from a height.the time taken by it to fall through succesive distances of 1m each will be

Grade:12th pass

1 Answers

Vishal
6 years ago
mgh = 1/2mv^2
gh = (v^2)/2
2*9.81*1=V^2
v = 4.42m/s
BY THIS you will get the answer.
P.E = K.E
PE = mgh and KE = 1/2 MV^2
hope you will understood.